OpenStudy (anonymous):

Does anyone know how to write a two colums proof If m 2 m 3 then m 3 m 2? I logged <2=<3, then <3=<2 under statement and reason is given, but after that I'm lost in the wind. brain stuck

OpenStudy (dumbcow):

not sure i understand the question. does m2 refer to the measure of an angle?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

if its given that m2 = m3 then yes you can say m3 = m2

OpenStudy (anonymous):

How do I complete the proof using previous attahcment

OpenStudy (dumbcow):

I think it would be the following: Given AB = VT and XB = XT If AX = AB - XB, and VX = VT - XT Then AX = VT - XT = VX

OpenStudy (dumbcow):

i know thats not in the correct form but thats the logic

OpenStudy (anonymous):

yes, I understand that proof, but how about the first one- If m< 2 m< 3 then m< 3 m< 2

OpenStudy (dumbcow):

i dont know it seems pretty straightforward, what is there to prove. m3 = m2 thus m2=m3. maybe check out a proof of vertical angles

OpenStudy (dumbcow):

sorry

OpenStudy (anonymous):

Thanks for your help